OpenStudy (gabrielah96):

solve for x. e^(1+2x)=5

OpenStudy (lochana):

\[\log_{e} e ^{1+2x} = \log_{e}5\] \[ (1+ 2x)\log_{e}e = \log_{e}5\]

OpenStudy (lochana):

\[(1+2x) = \log_{e}5\] \[x =\frac{ \log_{e}5 -1}{ 2 }\]

OpenStudy (lochana):

@gabrielah96 if you have any question about my post. feel free to ask:)

OpenStudy (gabrielah96):

thank you!!